# Reinforcement Learning in Vision: A Survey

### v2

@@ 4 Metrics and Benchmarks
Evaluating visual reinforcement learning (RL) with large models requires both traditional RL metrics (Jordan et al., 2020; Agarwal et al., 2021) and new ones designed for complex, open-ended outputs. Metrics like cumulative return and task success rate are still useful especially in tasks involving agents and environments but they are increasingly combined with preference-based evaluations. In this section, we discuss metrics and benchmarks across four major application areas of large-model RL.
@@ 4.1 Formalizing Metric Granularity
Task Metric $\mathcal{M}_{\mathrm{set}}$ (Set-level). As illustrated in Fig. 6 (a), set-level metrics evaluate the generative policy $\pi_{\theta}$ over the full prompt set $\mathcal{P}_{\text{test}}$, by comparing the distribution of generated outputs $\mathcal{Y}_{\text{gen}}=\{\pi_{\theta}(\cdot\mid p_{i})\}_{i=1}^{D}$ to a reference set of ground-truth outputs $\mathcal{Y}_{\text{gt}}=\{y^{i}_{\text{gt}}\}_{i=1}^{D}$. When the evaluation function operates without access to ground-truth outputs such as CLIPScore (Hessel et al., 2021), we define:
@@
where $\mathcal{M}_{\mathrm{samp}}(y,p)$ denotes a sample-level reward function applied to each generated output. By contrast, many classical vision metrics do rely on reference outputs such as IoU (Rezatofighi et al., 2019). For these cases, set-level evaluation is defined as:
@@
Reward/Advantage metric $\mathcal{M}_{\mathrm{samp}}$ (Sample-level). As illustrated in Fig. 6 (b), reward and advantage metrics $\mathcal{M}_{\mathrm{samp}}$ operate at the granularity of individual input-output pairs, forming the backbone of reinforcement learning in generative settings. Given a prompt $p_{i}$, the policy $\pi_{\theta}$ generates a sample $y^{i}_{\text{gen}}$, which is then scored by $\mathcal{M}_{\mathrm{samp}}(y^{i}_{\text{gen}},p_{i})$ to compute a scalar reward or advantage signal. This feedback is used to optimize the policy via reinforcement learning (e.g.,  PPO (Schulman et al., 2017), DPO (Rafailov et al., 2023)). In preference-based learning, the sample-level metric is often learned from human or GPT-4 comparisons (Ouyang et al., 2022; Sun et al., 2023), or automatically derived via scoring models like CLIPScore (Hessel et al., 2021), or ImageReward (Xu et al., 2023). When rewards are reference-dependent (e.g., using PSNR (Wang et al., 2004) or IoU (Rezatofighi et al., 2019)), $\mathcal{M}_{\mathrm{samp}}$ compares the generated output $y^{i}_{\text{gen}}$ to a ground-truth output $y^{i}_{\text{gt}}$. Formally, the reward signal can be expressed as:
@@ 4.2 Evaluation of Multi-Modal Large Language Models
Task Metric. As summarized in Table 2, MLLM are first judged on external, RL-free benchmarks. General reasoning suites such as MME (Fu et al., 2023), SEED-Bench (Li et al., 2024a) and MMBench (Liu et al., 2024b) measure factual QA, commonsense and multi-step chain-of-thought across images. Domain-specific subsets probe OCR (TextVQA (Singh et al., 2019)), mathematics (MathVista (Lu et al., 2023)), documents (ChartQA (Lu et al., 2023)) and multilingual grounding (CMMMU (Zhang et al., 2024a)).
Reward Metric. During training, each generated answer is scored with a sample-level reward $\mathcal{M}_{\mathrm{samp}}$. Three sources dominate current practice. (i) Human-preference rewards are learned from large RLHF corpora  e.g., InstructGPT (Ouyang et al., 2022) and LLaVA-RLHF (Sun et al., 2023), and give dense feedback that closely matches user judgements. (ii) Verifiable rewards arise when a sub-task admits deterministic checks, such as unit-test pass rate in CodeRL (Le et al., 2022) or symbolic exactness in DeepSeekMath (Shao et al., 2024b); they are noise-free but limited in scope. (iii) Model-preference rewards replace humans with a stronger frozen critic, e.g., CriticGPT (McAleese et al., 2024), delivering scalable but potentially biased supervision. The chosen reward is converted to advantages $\hat{A}_{i,t}$ and optimized via PPO, GRPO or DPO.
Model-State Metric. Beyond external scores, practitioners track light-weight diagnostics $\mathcal{M}_{\mathrm{state}}^{t}$ throughout RL updates. Two lightweight diagnostics are widely adopted: (i) Output length drift, the deviation of answer length from the SFT baseline large drift foreshadows verbosity or repetition (Rafailov et al., 2023); (ii) KL divergence between the current policy $\pi_{\theta_{t}}$ and frozen SFT reference $\pi_{\theta_{0}}$, as used in InstructGPT (Ouyang et al., 2022).
@@ 4.3 Evaluation of Visual Generation Models
Task Metric. As listed in Table 2, final quality is judged on standard, RL–free benchmarks that target complementary axes. Image fidelity & diversity is measured by FID and Inception Score, while pixel-level reconstruction tasks (super-resolution, inpainting) use PSNR or SSIM. For prompt alignment, CLIP Score and Fréchet CLIP Distance quantify semantic correctness; video models additionally report FVD or Video IS to capture temporal coherence.
Reward Metric. During RL fine-tuning, each generated image or video receives a sample-level reward $\mathcal{M}_{\mathrm{samp}}$. Human-preference rewards, e.g., ImageReward (Xu et al., 2023) and HPS (Wu et al., 2023b) supply dense signals that correlate well with aesthetic appeal. When a deterministic checker exists, authors turn to verifiable rewards: MotionPrompt (Nam et al., 2025) and DSPO (Cai et al., 2025b) use optical-flow, object masks that can be evaluated without humans. A third route relies on model preference rewards, where a stronger frozen critic (e.g., VideoPrefer (Wu et al., 2024) or PARM (Zhang et al., 2025f)) scores samples, enabling scalable DPO/PPO training.
Model-State Metric. Two light diagnostics track training stability. (i) Denoising trajectory statistics: VARD (Dai et al., 2025) and Inversion DPO (Li et al., 2025f) record per-step noise predictions or DDIM traces; pathological spikes reveal early collapse. (ii) KL divergence between the current diffusion policy and its frozen base ($\pi_{\theta_{t}}\!\parallel\!\pi_{\theta_{0}}$), popularized by DDPO (Black et al., 2023) and reused in VARD.
@@ 4.4 Evaluation of Unified Models
Task Metric. Two benchmark families are widely adopted. Generation-oriented suites such as GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024b), and ParaPrompts (Wu et al., 2025d) focus on prompt-faithful generation, testing multi-object composition, style control, and long-caption adherence. Conversely, understanding-oriented benchmarks (MME (Fu et al., 2023), POPE (Li et al., 2023)) measure grounding, reasoning and hallucination detection from the same backbone.
Reward Metric. Recent work explores two design philosophies for training signals. Unified rewardse.g.,   UniRL (Mao et al., 2025), CoRL (Jiang et al., 2025b)) blend multiple objectives textual correctness, CLIP alignment, aesthetic quality—into a single scalar that drives one shared policy across tasks. In contrast, task-specific rewards keep the generator and understanding heads separate, applying RL only to the generation branch as in Vargpt-v1.1 (Zhuang et al., 2025) or Emu3 (Wang et al., 2024b). The former promotes cross-modal transfer, while the latter preserves the stability of perception modules.
Model-State Metric. Unified models additionally track fine-grained diagnostics during RL. UniRL (Mao et al., 2025) proposes a generation–understanding imbalance score the absolute gap between batch-level rewards on the two task families to prevent one modality from dominating the update. HermesFlow (Yang et al., 2025c) monitors the KL divergence between the current shared policy $\pi_{\theta_{t}}$ and its supervised baseline $\pi_{\theta_{0}}$ on both generation and understanding prompts, serving as an early-warning signal for policy collapse. These state-level curves $\mathcal{M}_{\mathrm{state}}^{t}$ allow practitioners to stop or re-weight training before external task scores degrade.
@@ 4.5 Evaluation of Vision Language Action Models
Reward Metric. For reward modeling, most offline RL methods borrow the metric from task metric like IoU, while come to the online environment, due to the sparsity of task success rate, which present significant challenges for end-to-end multi-turn RL training (Lu et al., 2025a) i.e., lack of training efficency, lack of informativeness, step-level reward is proposed to address this issue, such as developing a reward or critic models (Bai et al., 2024; Chae et al., 2024).
@@
Table 3: Public benchmarks of MLLM most commonly used in visual RL. Only benchmarks relevant to visual reinforcement learning are included (RL-focused training & evaluation); task-specific benchmarks, such as MME (Fu et al., 2023), are excluded from consideration. ‘Tr’ and ‘Te’ refer to the ‘Train’ and ‘Test’, respectively.
@@
SEED-Bench-R1 (Chen et al., 2025c) Mar 2025 Tr&Te Video-QA pairs with human-preference reward model
Video-Holmes (Cheng et al., 2025) May 2025 Te Human-ranked T2I pairs from diverse generation models
VisuLogic (Xu et al., 2025a) Apr 2025 Tr&Te Visual-reasoning QA set; exact-match reward enables RL fine-tuning
R1-ShareVL (Yao et al., 2025) May 2025 Tr 52 k MM-Eureka subset for Share-GRPO RL training
Ego-R1 (Tian et al., 2025) Jun 2025 Tr&Te 25 k CoTT egocentric traces enabling RL training for ultra-long video reasoning
Long-RL (Chen et al., 2025f) Jul 2025 Tr&Te 104 K long-video QA pairs (GRPO accuracy / format reward)
VisCOT (Shao et al., 2024a) Mar 2024 Tr&Te 438 k visual chain-of-thought traces with step-wise rewards for RL
MMK12 (Meng et al., 2025) Mar 2025 Tr&Te 15.6 K multimodal math problems (rule-based accuracy / format rewards)
Time-R1 (Wang et al., 2025h) Mar 2025 Tr&Te 2.5 K TimeRFT grounding spans (IoU reward)
VSI-Bench (Yang et al., 2024) Dec 2024 Te Spatial QA benchmark offering RL exact-match reward
MME-Reasoning (Yuan et al., 2025a) May 2025 Te Logic QA benchmark
K12-2M (Wang et al., 2025d) May 2025 Tr 2 M multimodal math pairs for RL training
MathVista (Lu et al., 2024) Oct 2023 Te Visual-math QA benchmark; test-only RL exact-match reward
MathVerse (Zhang et al., 2024c) Mar 2024 Te Diagram-math QA benchmark; test-only RL exact-match reward
EMMA (Hao et al., 2025) Jan 2025 Te Robust multimodal reasoning benchmark challenging current MLLMs
WeMath (Qiao et al., 2024) Jul 2024 Te Multimodal math reasoning benchmark with diagrams
DynaMATH (Zou et al., 2024) Oct 2024 Te Dynamic visual math reasoning robustness benchmark for VLMs
MM-IQ (Cai et al., 2025a) Feb 2025 Te Human-like multimodal abstraction and reasoning benchmark
OlympiadBench (He et al., 2024a) Feb 2024 Te Bilingual multimodal Olympiad-level scientific reasoning benchmark
ZeroBench (Roberts et al., 2025) Feb 2025 Te near-impossible visual reasoning stress test for LMMs
MMMU-Pro (Yue et al., 2024) Sep 2024 Te robust multidisciplinary multimodal understanding benchmark
MME-CoT (Jiang et al., 2025a) Feb 2025 Te Multimodal chain-of-thought reasoning benchmark dataset
MMIR (Yan et al., 2025) Feb 2025 Te Multimodal inconsistency reasoning benchmark dataset
SpatialEval (Wang et al., 2024a) Jun 2024 Te Synthetic spatial reasoning benchmark for VLMs
MMReason (Wang et al., 2024a) Jun 2025 Te Synthetic spatial reasoning benchmark for VLMs
IntentBench (Yang et al., 2025d) Jun 2025 Te Omnimodal evaluation demands unified audio-visual comprehension
Table 4: Public benchmarks of Visual Generation (image/video/3D) most commonly used in visual RL. Only benchmarks relevant to visual reinforcement learning are included (RL-focused training & evaluation). ‘Tr’ and ‘Te’ refer to the ‘Train’ and ‘Test’, respectively.

### v3

@@ 4 Metrics and Benchmarks
Evaluating visual reinforcement learning (RL) with large models requires both traditional RL metrics [jordan2020evaluating, agarwal2021deep] and new ones designed for complex, open-ended outputs. Metrics like cumulative return and task success rate are still useful especially in tasks involving agents and environments but they are increasingly combined with preference-based evaluations. In this section, we discuss metrics and benchmarks across four major application areas of large-model RL.
@@ 4.1 Formalizing Metric Granularity
Task Metric $\mathcal{M}_{\mathrm{set}}$ (Set-level). As illustrated in Fig. 6 (a), set-level metrics evaluate the generative policy $\pi_{\theta}$ over the full prompt set $\mathcal{P}_{\text{test}}$, by comparing the distribution of generated outputs $\mathcal{Y}_{\text{gen}}=\{\pi_{\theta}(\cdot\mid p_{i})\}_{i=1}^{D}$ to a reference set of ground-truth outputs $\mathcal{Y}_{\text{gt}}=\{y^{i}_{\text{gt}}\}_{i=1}^{D}$. When the evaluation function operates without access to ground-truth outputs such as CLIPScore [hessel2021clipscore], we define:
@@
where $\mathcal{M}_{\mathrm{samp}}(y,p)$ denotes a sample-level reward function applied to each generated output. By contrast, many classical vision metrics do rely on reference outputs such as IoU [rezatofighi2019generalized]. For these cases, set-level evaluation is defined as:
@@
Reward/Advantage metric $\mathcal{M}_{\mathrm{samp}}$ (Sample-level). As illustrated in Fig. 6 (b), reward and advantage metrics $\mathcal{M}_{\mathrm{samp}}$ operate at the granularity of individual input-output pairs, forming the backbone of reinforcement learning in generative settings. Given a prompt $p_{i}$, the policy $\pi_{\theta}$ generates a sample $y^{i}_{\text{gen}}$, which is then scored by $\mathcal{M}_{\mathrm{samp}}(y^{i}_{\text{gen}},p_{i})$ to compute a scalar reward or advantage signal. This feedback is used to optimize the policy via reinforcement learning (e.g.,  PPO [schulman2017proximal], DPO [rafailov2023direct]). In preference-based learning, the sample-level metric is often learned from human or GPT-4 comparisons [ouyang2022training, sun2023aligning], or automatically derived via scoring models like CLIPScore [hessel2021clipscore], or ImageReward [xu2023imagereward]. When rewards are reference-dependent (e.g., using PSNR [wang2004image] or IoU [rezatofighi2019generalized]), $\mathcal{M}_{\mathrm{samp}}$ compares the generated output $y^{i}_{\text{gen}}$ to a ground-truth output $y^{i}_{\text{gt}}$. Formally, the reward signal can be expressed as:
@@ 4.2 Evaluation of Multi-Modal Large Language Models
Task Metric. As summarized in Table II, MLLM are first judged on external, RL-free benchmarks. General reasoning suites such as MME [fu2023mme], SEED-Bench [li2024seed] and MMBench [liu2024mmbench] measure factual QA, commonsense and multi-step chain-of-thought across images. Domain-specific subsets probe OCR (TextVQA [singh2019towards]), mathematics (MathVista [lu2023mathvista]), documents (ChartQA [lu2023mathvista]) and multilingual grounding (CMMMU [zhang2024cmmmu]).
Reward Metric. During training, each generated answer is scored with a sample-level reward $\mathcal{M}_{\mathrm{samp}}$. Three sources dominate current practice. (i) Human-preference rewards are learned from large RLHF corpora  e.g., InstructGPT [ouyang2022training] and LLaVA-RLHF [sun2023aligning], and give dense feedback that closely matches user judgements. (ii) Verifiable rewards arise when a sub-task admits deterministic checks, such as unit-test pass rate in CodeRL [le2022coderl] or symbolic exactness in DeepSeekMath [shao2024deepseekmath]; they are noise-free but limited in scope. (iii) Model-preference rewards replace humans with a stronger frozen critic, e.g., CriticGPT [mcaleese2024llm], delivering scalable but potentially biased supervision. The chosen reward is converted to advantages $\hat{A}_{i,t}$ and optimized via PPO, GRPO or DPO.
Model-State Metric. Beyond external scores, practitioners track light-weight diagnostics $\mathcal{M}_{\mathrm{state}}^{t}$ throughout RL updates. Two lightweight diagnostics are widely adopted: (i) Output length drift, the deviation of answer length from the SFT baseline large drift foreshadows verbosity or repetition [rafailov2023direct]; (ii) KL divergence between the current policy $\pi_{\theta_{t}}$ and frozen SFT reference $\pi_{\theta_{0}}$, as used in InstructGPT [ouyang2022training].
@@ 4.3 Evaluation of Visual Generation Models
Task Metric. As listed in Table II, final quality is judged on standard, RL–free benchmarks that target complementary axes. Image fidelity & diversity is measured by FID and Inception Score, while pixel-level reconstruction tasks (super-resolution, inpainting) use PSNR or SSIM. For prompt alignment, CLIP Score and Fréchet CLIP Distance quantify semantic correctness; video models additionally report FVD or Video IS to capture temporal coherence.
Reward Metric. During RL fine-tuning, each generated image or video receives a sample-level reward $\mathcal{M}_{\mathrm{samp}}$. Human-preference rewards, e.g., ImageReward [xu2023imagereward] and HPS [wu2023human] supply dense signals that correlate well with aesthetic appeal. When a deterministic checker exists, authors turn to verifiable rewards: MotionPrompt [nam2025optical] and DSPO [cai2025dspo] use optical-flow, object masks that can be evaluated without humans. A third route relies on model preference rewards, where a stronger frozen critic (e.g., VideoPrefer [wu2024boosting] or PARM [zhang2025let]) scores samples, enabling scalable DPO/PPO training.
Model-State Metric. Two light diagnostics track training stability. (i) Denoising trajectory statistics: VARD [dai2025vard] and Inversion DPO [li2025inversion] record per-step noise predictions or DDIM traces; pathological spikes reveal early collapse. (ii) KL divergence between the current diffusion policy and its frozen base ($\pi_{\theta_{t}}\!\parallel\!\pi_{\theta_{0}}$), popularized by DDPO [black2023ddpo] and reused in VARD.
@@ 4.4 Evaluation of Unified Models
Task Metric. Two benchmark families are widely adopted. Generation-oriented suites such as GenEval [ghosh2023geneval], DPG-Bench [hu2024ella], and ParaPrompts [wu2025paragraph] focus on prompt-faithful generation, testing multi-object composition, style control, and long-caption adherence. Conversely, understanding-oriented benchmarks (MME [fu2023mme], POPE [li2023evaluating]) measure grounding, reasoning and hallucination detection from the same backbone.
Reward Metric. Recent work explores two design philosophies for training signals. Unified rewardse.g.,   UniRL [mao2025unirl], CoRL [jiang2025co]) blend multiple objectives textual correctness, CLIP alignment, aesthetic quality—into a single scalar that drives one shared policy across tasks. In contrast, task-specific rewards keep the generator and understanding heads separate, applying RL only to the generation branch as in Vargpt-v1.1 [zhuang2025vargpt] or Emu3 [wang2024emu3]. The former promotes cross-modal transfer, while the latter preserves the stability of perception modules.
Model-State Metric. Unified models additionally track fine-grained diagnostics during RL. UniRL [mao2025unirl] proposes a generation–understanding imbalance score the absolute gap between batch-level rewards on the two task families to prevent one modality from dominating the update. HermesFlow [yang2025hermesflow] monitors the KL divergence between the current shared policy $\pi_{\theta_{t}}$ and its supervised baseline $\pi_{\theta_{0}}$ on both generation and understanding prompts, serving as an early-warning signal for policy collapse. These state-level curves $\mathcal{M}_{\mathrm{state}}^{t}$ allow practitioners to stop or re-weight training before external task scores degrade.
@@ 4.5 Evaluation of Vision Language Action Models
Reward Metric. For reward modeling, most offline RL methods borrow the metric from task metric like IoU, while come to the online environment, due to the sparsity of task success rate, which present significant challenges for end-to-end multi-turn RL training [arpo] i.e., lack of training efficency, lack of informativeness, step-level reward is proposed to address this issue, such as developing a reward or critic models [digirl, webworld].
@@
TABLE III: Public benchmarks of MLLM most commonly used in visual RL. Only benchmarks relevant to visual reinforcement learning are included (RL-focused training & evaluation); task-specific benchmarks, such as MME [fu2023mme], are excluded from consideration. ‘Tr’ and ‘Te’ refer to the ‘Train’ and ‘Test’, respectively.
Task family Benchmark Date Tr/Te Description (benchmark info. and RL reward signal)
Multimodal LLMs / VLMs SEED-Bench-R1 [chen2025exploring] Mar 2025 Tr&Te Video-QA pairs with human-preference reward model (used in [chen2025exploring])
Video-Holmes [cheng2025video] May 2025 Te Human-ranked T2I pairs from diverse generation models (used in [cheng2025video])
VisuLogic [xu2025visulogic] Apr 2025 Tr&Te Visual-reasoning QA set; exact-match reward enables RL fine-tuning (used in [xu2025visulogic])
R1-ShareVL [yao2025r1sharevl] May 2025 Tr 52 k MM-Eureka subset for Share-GRPO RL training (used in [yao2025r1sharevl])
Ego-R1 [tian2025egor1chainoftoolthoughtultralongegocentric] Jun 2025 Tr&Te 25 k CoTT egocentric traces enabling RL training for ultra-long video reasoning (used in [tian2025egor1chainoftoolthoughtultralongegocentric])
Long-RL [chen2025scaling] Jul 2025 Tr&Te 104 K long-video QA pairs (GRPO accuracy / format reward) (used in [chen2025scaling])
VisCOT [shao2024visual] Mar 2024 Tr&Te 438 k visual chain-of-thought traces with step-wise rewards for RL (used in [shao2024visual])
MMK12 [meng2025mm] Mar 2025 Tr&Te 15.6 K multimodal math problems (rule-based accuracy / format rewards) (used in [meng2025mm])
Time-R1 [wang2025timer1] Mar 2025 Tr&Te 2.5 K TimeRFT grounding spans (IoU reward) (used in [wang2025timer1])
VSI-Bench [yang2024think] Dec 2024 Te Spatial QA benchmark offering RL exact-match reward (used in [yang2024think])
MME-Reasoning [yuan2025mme] May 2025 Te Logic QA benchmark (used in [yuan2025mme])
K12-2M [wang2025mathcodervl] May 2025 Tr 2 M multimodal math pairs for RL training (used in [wang2025mathcodervl])
MathVista [lu2024mathvista] Oct 2023 Te Visual-math QA benchmark; test-only RL exact-match reward (used in [lu2024mathvista])
MathVerse [zhang2024mathverse] Mar 2024 Te Diagram-math QA benchmark; test-only RL exact-match reward (used in [zhang2024mathverse])
EMMA [hao2025can] Jan 2025 Te Robust multimodal reasoning benchmark challenging current MLLMs (used in [hao2025can])
WeMath [qiao2024we] Jul 2024 Te Multimodal math reasoning benchmark with diagrams (used in [qiao2024we])
DynaMATH [zou2024dynamic] Oct 2024 Te Dynamic visual math reasoning robustness benchmark for VLMs (used in [zou2024dynamic])
MM-IQ [cai2025mm] Feb 2025 Te Human-like multimodal abstraction and reasoning benchmark (used in [cai2025mm])
OlympiadBench [he2024olympiadbench] Feb 2024 Te Bilingual multimodal Olympiad-level scientific reasoning benchmark (used in [he2024olympiadbench])
ZeroBench [roberts2025zerobench] Feb 2025 Te near-impossible visual reasoning stress test for LMMs (used in [roberts2025zerobench])
MMMU-Pro [yue2024mmmu] Sep 2024 Te robust multidisciplinary multimodal understanding benchmark (used in [yue2024mmmu])
MME-CoT [jiang2025mme] Feb 2025 Te Multimodal chain-of-thought reasoning benchmark dataset (used in [jiang2025mme])
MMIR [yan2025multimodalinconsistencyreasoningmmir] Feb 2025 Te Multimodal inconsistency reasoning benchmark dataset (used in [yan2025multimodalinconsistencyreasoningmmir])
SpatialEval [wang2024spatial] Jun 2024 Te Synthetic spatial reasoning benchmark for VLMs (used in [wang2024spatial])
MMReason [wang2024spatial] Jun 2025 Te Synthetic spatial reasoning benchmark for VLMs (used in [wang2024spatial])
IntentBench [yang2025humanomniv2] Jun 2025 Te Omnimodal evaluation demands unified audio-visual comprehension (used in [yang2025humanomniv2])
Visual Generation (image/video/3D) ImageReward [xu2023imagereward] Apr 2023 Tr&Te Human-ranked pairs for T2I; reward model outputs scalar score (used in [wu2025reprompt, xiao2024diffppo, zhou2025multimodalllmscustomizedreward])
HPS [wu2023human] Mar 2023 Tr&Te Human-ranked T2I pairs from diverse generation models (used in [wu2023human])
HPS V2 [wu2023humanv2] Jun 2023 Tr&Te Human-ranked T2I pairs across diverse prompts, preference-based reward (used in [duan2025got, gu2024diffusion])
Pick-a-Pic [kirstain2023pick] May 2023 Tr&Te User preferences for pairwise rankings over T2I generations (used in [wallace2024diffusion, li2024aligning, gu2024diffusion, lee2024parrot])
VideoReward [liu2025improving] Jan 2025 Te Human-ranked video pairs across quality, motion, and text alignment (used in [liu2025improving])
T2I-CompBench [huang2023t2i] Jul 2023 Tr&Te Compositional text-to-image dataset covering attributes and object relations. (used in [gupta2025simple, duan2025got])
StarVector [rodriguez2025starvector] Dec 2023 Tr&Te SVG Code Generation Data, Match Reward (used in [rodriguez2025rendering])
AnimeReward [zhu2025aligning] Apr 2025 Tr Multi‑dimensional (e.g., character consisten) human preference anime videos (used in [zhu2025aligning])
VideoPrefer [wu2024boosting] Dec 2024 Tr MLLM‑annotated 135K video preference pairs (used in [wu2024boosting])
Vision–Language Action Agents GUI-R1-3K [luo2025gui] Apr 2025 Tr&Te GUI trajectories spanning Windows, Linux, macOS, Android, and Web platforms (used in [luo2025gui])
SE-GUI-3k [yuan2025enhancing] May 2025 Tr 3,018 examples (desktop / web / mobile) with instruction and bounding box (used in [yuan2025enhancing])
UI-R1 [lu2025ui] May 2025 Tr 136 mobile GUI tasks covering 5 action types (click, scroll, swipe, text-input) (used in [yuan2025enhancing])
CAGUI [zhang2025agentcpm] Jun 2025 Te 55 K trajectories from 30 Chinese Android apps, 8 domains (used in [yuan2025enhancing])
Mobile-R1 [gu2025mobile] Jun 2025 Tr&Te More than 500 online task trajectories from 28 Chinese apps (used in [gu2025mobile])
Mind2web [deng2023mind2web] Jun 2023 Tr&Te 2 k tasks on 137 real websites; success/fail reward for RL (used in [deng2023mind2web, tang2025lpo])
AITZ [zhang2024android] Jun 2023 Tr&Te 18,643 Android screen–action pairs with CoAT reasoning (used in [zhang2024android])
Omniac [kapoor2024omniact] Feb 2024 Tr&Te Desktop + web  9.8 k scripted tasks (used in [kapoor2024omniact])
GUICours [chen2024guicourse] Jun 2024 Tr&Te GUIEnv/Act/Chat datasets (10 M OCR + 67 k navigation) (used in [chen2024guicourse])
Habitat 3.0 [puig2023habitat] Oct 2023 Tr&Te Interactive embodied-AI scenes with humans and robots (used in [puig2023habitat])
VLN-CE [krantz2020beyond] Apr 2020 Tr&Te Continuous embodied navigation dataset with language instructions (used in [krantz2020beyond])
Rlbench [james2020rlbench] Apr 2020 Tr&Te Multi-task simulated robot manipulation benchmark dataset (used in [james2020rlbench])
RoboCasa [nasiriany2024robocasa] Jun 2024 Tr Large-scale kitchen-task simulation for generalist robotics (used in [nasiriany2024robocasa])
LIBERO [liu2023libero] Jun 2023 Tr Lifelong robot learning benchmark with 100 manipulation tasks (used in [liu2023libero])
VLABench [zhang2024vlabench] Dec 2024 Tr&Te Long-horizon language-conditioned manipulation benchmark for robots (used in [zhang2024vlabench])
TABLE IV: Public benchmarks of Visual Generation (image/video/3D) most commonly used in visual RL. Only benchmarks relevant to visual reinforcement learning are included (RL-focused training & evaluation). ‘Tr’ and ‘Te’ refer to the ‘Train’ and ‘Test’, respectively.
@@
ImageReward [xu2023imagereward] Apr 2023 Tr&Te Human-ranked pairs for T2I; reward model outputs scalar score (used in [wu2025reprompt, xiao2024diffppo, zhou2025multimodalllmscustomizedreward])
HPS [wu2023human] Mar 2023 Tr&Te Human-ranked T2I pairs from diverse generation models (used in [wu2023human])
HPS V2 [wu2023humanv2] Jun 2023 Tr&Te Human-ranked T2I pairs across diverse prompts, preference-based reward (used in [duan2025got, gu2024diffusion])
Pick-a-Pic [kirstain2023pick] May 2023 Tr&Te User preferences for pairwise rankings over T2I generations (used in [wallace2024diffusion, li2024aligning, gu2024diffusion, lee2024parrot])
VideoReward [liu2025improving] Jan 2025 Te Human-ranked video pairs across quality, motion, and text alignment (used in [liu2025improving])
T2I-CompBench [huang2023t2i] Jul 2023 Tr&Te Compositional text-to-image dataset covering attributes and object relations. (used in [gupta2025simple, duan2025got])
StarVector [rodriguez2025starvector] Dec 2023 Tr&Te SVG Code Generation Data, Match Reward (used in [rodriguez2025rendering])
AnimeReward [zhu2025aligning] Apr 2025 Tr Multi‑dimensional (e.g., character consisten) human preference anime videos (used in [zhu2025aligning])
VideoPrefer [wu2024boosting] Dec 2024 Tr MLLM‑annotated 135K video preference pairs (used in [wu2024boosting])
TABLE V: Public benchmarks of VLAs most commonly used in visual RL. Only benchmarks relevant to visual reinforcement learning are included (RL-focused training & evaluation). ‘Tr’ and ‘Te’ refer to the ‘Train’ and ‘Test’, respectively.